OpenStudy (vera_ewing):

What would you expect to see when silver is placed in hydrochloric acid? A. Bubbles of chlorine gas would form. B. The silver metal would disappear. C. Silver ions would become solid silver. D. No reaction would happen.

OpenStudy (vera_ewing):

@nechirwan

OpenStudy (nechirwan):

Ag cannot displace H in HCl ; HCl is not strong enough acid but it reacts with concentrated nitric acid, as well as with concentrated sulfuric acid. like this: HNO3 + Ag = AgNO3 + H2O + NO H2SO4 + Ag = Ag2SO4 + H2O + SO2

OpenStudy (vera_ewing):

@nechirwan so what's the answer?

OpenStudy (nechirwan):

D. No reaction would happen.

OpenStudy (vera_ewing):

thanks! :)